40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Mar 2 21:16:32 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
    Reply  Tue Mar 3 19:26:44 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
       Reply  Tue Mar 3 23:59:26 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
          Reply  Wed Mar 4 03:37:51 2009, Yoichi, Update, Locking, Low-gain High-gain PD switching may not be working well 
Message ID: 1346     Entry time: Mon Mar 2 21:16:32 2009     Reply to this: 1350
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: Low-gain High-gain PD switching may not be working well 
Osamu, Yoichi

This afternoon, I run the locking script while doing calculations for the upgrade.
The IFO lost lock even at lower arm powers (around 6) if it was operated for a while (~ 5min).
It seemed as if there were some intermittent glitches (seismic? laser?) causing the lock losses.
We also saw once the TRX and TRY signals saturated at around arm power = 11 when there was a large fluctuation in the arm power.
Osamu suggested that it looked like the high-gain to low-gain PD switching was not working.

I won't come tonight as I may have caught a cold, but if someone comes tonight, it is worth checking the PD switching.
ELOG V3.1.3-